OpenStudy (anonymous):

Sally currently has an account balance of $991.44. Her initial deposit on the account was $540 and it earned 4.4% simple interest. How long has Sally held the account? 10 years 42 years 23 years 19 years

OpenStudy (nubeer):

991. 44 = 540 *4.4% * t t = (991.44/540*4.4%) find t u will get the answer

OpenStudy (anonymous):

thanks! :)

OpenStudy (nubeer):

you're welcome :)